### v1

# STATE ToxiCN: A Benchmark for Span-level Target-Aware Toxicity Extraction in Chinese Hate Speech Detection

Abstract
The proliferation of hate speech has caused significant harm to society. The intensity and directionality of hate are closely tied to the target and argument it is associated with. However, research on hate speech detection in Chinese has lagged behind, and existing datasets lack span-level fine-grained annotations. Furthermore, the lack of research on Chinese hateful slang poses a significant challenge. In this paper, we provide a solution for fine-grained detection of Chinese hate speech. First, we construct a dataset containing Target-Argument-Hateful-Group quadruples (STATE ToxiCN), which is the first span-level Chinese hate speech dataset. Secondly, we evaluate the span-level hate speech detection performance of existing models using STATE ToxiCN. Finally, we conduct the first study on Chinese hateful slang and evaluate the ability of LLMs to detect such expressions. Our work contributes valuable resources and insights to advance span-level hate speech detection in Chinese. [^1] Code and datasets are publicly available at https://github.com/shenmeyemeifashengguo/STATE-ToxiCN.
Disclaimer: The samples presented by this paper may be considered offensive or vulgar.
1 Introduction
With the popularity of social media, user-generated content has experienced explosive growth, and hate speech has also flourished. Hate speech refers to harmful statements that express hatred or incite harm against specific groups or individuals based on race, religion, gender, region, sexual orientation, physical characteristics, and other factors Bilewicz and Soral (2020). These statements are often driven by group hysteria, personal experiences, and social trends, and are characterized by compulsiveness, bullying, and incitement. Approximately 941 million people, or 12% of the global population, speak Mandarin Chinese as their first language Eberhard et al. (2024). The issue of effectively detecting and managing Chinese hate speech has become a topic of widespread concern.
The intensity and directionality of hate speech are closely tied to the target and argument it is associated with Cowan and Hodge (1996). In recent years, research on hate speech detection in English has gradually shifted from post-level Davidson et al. (2017); Founta et al. (2018) to span-level Pavlopoulos et al. (2021); Mathew et al. (2021); Zampieri et al. (2023). However, in the Chinese domain, span-level hate speech detection remains unexplored, significantly limiting the development of Chinese hate speech detection. This task faces unique challenges within the Chinese linguistic environment. Firstly, Chinese syntax’s flexible word order makes identifying the target and argument ambiguous. For example, as shown in Exp.1 in Table 1, the use of inversion in Chinese disrupts the conventional subject-verb-object structure. Secondly, the absence of word delimiters in Chinese, such as the spaces used in English, complicates word segmentation. Furthermore, the scarcity of annotated resources exacerbates this issue, highlighting the urgent need for research in span-level Chinese hate speech detection.
Exp. Post Target Argument Hateful Group
1 你这头蠢驴,没人会喜欢。 蠢驴 non-hate non-hate
No one will ever like you, you idiot. you idiot
2 男同是艾滋高发群体。 男同 艾滋高发群体 hate LGBTQ, others
Gay people are a high-risk group for HIV. Gay people a high-risk group for HIV
3 默我是真的很讨厌。 默(黑犬) 讨厌 hate Racism
Silence, I really hate it. Slience(black dog) hate
Table 1: Examples of annotated posts from STATE ToxiCN dataset with corresponding annotations of Target-Argument-Hateful-Group quadruples.
The ideographic nature of Chinese, with its abundance of synonyms and near-synonyms, leads to highly varied hateful slang. Chinese hateful slang often evades model detection through various techniques such as homophonic substitutions, character splitting and merging, and historical allusions, which severely compromise the accuracy of existing models. For example, in Exp.3 from Table 1, the target uses a merge-word technique to evade detection. However, previous work has only provided words, lacking interpretable annotations. To address this gap, we collect and summarize commonly used hateful slang from real-world online forums, providing detailed annotations to construct the first annotated lexicon for Chinese hateful slang. This lexicon offers a reliable benchmark for Chinese hate speech detection and lays a foundation for improving model performance.
To address the lack of resources for span-level Chinese hate speech research, we introduce a Span-level Target-Aware Toxicity Extraction dataset (STATE ToxiCN), a novel dataset containing 8029 posts and 9533 quadruples addressing sexism, racism, regional bias, and anti-LGBTQ sentiments. Using this dataset, we evaluate the performance of LLMs in span-level Chinese hate speech detection. Specially, we annotate: 1) extraction of the targets and arguments from the post, 2) determination of whether each Target-Argument pair constitutes hate speech, and 3) classification of the groups for hateful Target-Argument pairs.
Moreover, we systematically summarize and annotate commonly used hateful slang from real-world online forums to address the challenges of identifying hateful slang. We compile a comprehensive annotated lexicon, labeling each hateful slang with its frequent groups and providing detailed explanations of its usage and contextual meaning. This resource not only helps to understand how hateful slang are used to disguise hate speech but also provides valuable annotated data for evaluating and improving the ability of LLMs to detect Chinese hateful slang. The main contributions of this work are summarized as follows:
  • We provide a span-level target-aware toxicity extraction dataset, containing 8k posts and 9.5k quadruples, filling the gap in span-level resources for Chinese hate speech detection.
  • We construct an annotated lexicon of commonly used hateful slang in online forums, including their frequent groups and contextual explanations as annotations.
  • We evaluate models on STATE ToxiCN, assessing their span-level performance and ability to detect hateful slang, highlighting key challenges and insights for improvement.
2 Related Work
Work Platforms Language #Posts Span Tar. Arg. Group Lex.
Davidson et al. (2017) Twitter English 24,802
Founta et al. (2018) Twitter English 80,000
Toxic Spans Pavlopoulos et al. (2021) Civil Comments English 10,629
HateXplain Mathew et al. (2021) Twitter, Gab English 20,148
TBO Zampieri et al. (2023) Twitter English 4,673
[…]
SWSR Jiang et al. (2022) Weibo Chinese 8,969
Cdial-Bias-Utt Zhou et al. (2022) Zhihu Chinese 13,394
Cdial-Bias-Ctx Zhou et al. (2022) Zhihu Chinese 15,013
ToxiCN Lu et al. (2023) Zhihu, Tieba Chinese 12,011
[…]
Table 2: Comparison of hate speech datasets based on Platforms, Language, #Posts, span-level annotations (Span), inclusion of Target (Tar.), Argument (Arg.), Group, and Lexicon information (Lex.).
Hate Speech Detection.
Hate speech detection is a critical task in NLP that has attracted considerable attention recently. Researchers have increasingly turned to pre-trained models to address this issue Caselli et al. (2020); Hanu and Unitary team (2020); Zhou et al. (2021); AlKhamissi et al. (2022); Ali et al. (2022). To facilitate progress in this field, several datasets tailored to hate speech detection have been developed Waseem and Hovy (2016); Davidson et al. (2017); Founta et al. (2018); Hartvigsen et al. (2022). Pavlopoulos et al. (2021) introduced span-level hate speech detection, while the TBO dataset advanced the field by pioneering the extraction of Target-Argument-Harmful triples Zampieri et al. (2023). However, Chinese hate speech detection remains significantly underdeveloped.
Chinese Hate Speech Dataset.
While some Chinese hate speech datasets exist, these efforts remain limited to the post-level. TOCP and TOCAB, derived from Taiwan’s PTT platform, focus on detecting profanity and abusive language Chung and Lin (2021). The Sina Weibo Sexism Review (SWSR) centers on identifying sexism Jiang et al. (2022). The Chinese Offensive Language Dataset (COLD) categorizes sentences into types such as individual attacks and anti-bias Deng et al. (2022). Zhou et al. (2022) introduce CDial-Bias, the first annotated dataset specifically addressing social bias in Chinese dialogues. Lu et al. (2023) present ToxiCN, a dataset encompassing both explicit and implicit toxic language samples.
While previous works offer quality corpora, span-level Chinese hate speech detection remains unexplored. STATE ToxiCN is the first such dataset, annotating Target-Argument-Hateful- Group quadruples within posts. Unlike previous lexicons, we provide interpretative annotations and targeted group labels, creating a comprehensive lexicon of Chinese hateful slang.
3 Dataset Construction
3.1 Overview
In this section, we introduce the construction process of the STATE ToxiCN dataset and the annotated lexicon of Chinese hateful slang. First, we describe the data sources and filtering procedures. Next, we detail the annotation process and the measures implemented to ensure high annotation quality. We then examine the Inter-Annotator Agreement (IAA) at different levels of granularity. Finally, we present relevant statistics for the STATE ToxiCN dataset.
3.2 Data Source and Filtering
Our dataset construction is based on the post-level Chinese hate speech dataset ToxiCN Lu et al. (2023). We develop a high-quality span-level Chinese hate speech dataset through data filtering and annotation processes. During the data filtering phase, we extract potential samples from the original dataset and systematically clean low-quality, incomplete, or irrelevant content. For instance, we remove meaningless text such as advertisements, random character combinations, and overly short (less than 5 characters) or overly long (more than 500 characters) text fragments.
Building on this, we conduct a comprehensive review of the hate speech annotations, particularly regarding the clarity of the target and argument components of hate speech. For example, in cases describing strongly discriminatory or biased content, we further examine whether the target was specific, removing ambiguous or undefined samples. In particular, for texts lacking a clear Target-Argument structure, we opt to exclude them to ensure the dataset could accommodate the requirements of span-level analysis.
3.3 Data Annotation
3.3.1 Annotation Guidelines
During the annotation process, we adhere to strict guidelines and implement multi-stage quality control to ensure the accuracy and consistency of the annotations, ultimately constructing a high-quality span-level Chinese hate speech dataset. First, we develop detailed annotation guidelines, including standards for extracting targets and arguments (Target-Argument Pair), criteria for determining hatefulness (Hateful), and methods for classifying groups (Group):
Target-Argument Pair
A combination of the target and the argument extracted from the post. A single post may contain more than one Target-Argument Pair.
Hateful
If the Target-Argument Pair explicitly or implicitly conveys harm towards the target or other groups, it is labeled as "hateful"; otherwise, it is labeled as "non-hate."
Group
Building on the Target-Argument Pair, this category identifies specific groups targeted by hateful expressions, with a single pair probably involving multiple groups.
During the construction of the annotated lexicon for Chinese hateful slang, we follow well-defined guidelines and employ systematic quality control to ensure reliability. This process include detailed criteria for identifying, categorizing, and annotating hateful slang terms, with a focus on their contextual usage and targeted implications:
Group
Each hateful slang term is categorized by the group it targets, such as sexism, racism, regional bias, anti-LGBTQ, or others. Some terms may target multiple groups.
Explanation
An explanation of hateful slang is provided, including its literal meaning, extended meanings, the reasons for hatred towards targeted groups, and common usage patterns.
3.3.2 Mitigating Bias.
To mitigate bias, we assemble annotators with diverse backgrounds, including differences in gender, age, ethnicity, region, and educational level. We believe this diversity helps reduce bias from a single perspective and enhances the reliability of the annotation results. All annotators possess linguistic expertise and have undergone systematic training. During the annotation process, we primarily employ regular cross-validation and expert arbitration to ensure the objectivity and consistency of annotation results. Additionally, we maintain an online document to record and update Chinese hateful slang identified in the posts. This annotated lexicon serves not only as a research resource but also helps to align annotators’ standards.
3.3.3 Annotation Procedure
Target Span Argument Span If Hateful Targeted Group
0.65 0.61 0.68 0.75
Table 3: Fleiss’ Kappa for different granularities.
Cross-validation and Expert Arbitration.
Each text is independently annotated by at least two annotators, who followed a unified annotation guideline to extract Target-Argument pairs, determine the hatefulness, and classify the groups. After the initial annotation phase, a portion of the samples is regularly selected for cross-validation, during which other annotators re-annotate these samples to ensure consistent understanding of the rules and standards across annotators. This approach allows us to identify and resolve potential biases or discrepancies in the annotations in a timely manner.
For disputed samples, where significant differences in annotation results are observed between annotators, the samples are submitted to an arbitration team composed of domain experts for review. The expert team analyze the disputed samples in depth, taking into account the textual context and the annotation guidelines, and determined the most reasonable annotation results. After annotation, we evaluate inter-annotator agreement (IAA) on STATE ToxiCN, with Fleiss’ kappa scores for each hierarchy detailed in Table 3.
Category Subcategory Count Percentage (%)
Groups Gender 1663 17.44
Race 1232 12.92
Region 1323 13.88
LGBTQ 628 6.59
Others 351 3.68
Multi-group 866 9.08
Hateful Hate 6063 63.60
Non-Hate 3470 36.40
Total - 9533 100.00
Table 4: Statistics of annotated posts from the STATE ToxiCN dataset, including Target, Argument, Group, and Hatefulness classifications.
Lexicon Annotation.
To ensure annotation quality, we maintain a shared online document for recording and updating detailed information on Chinese hateful slang. This document includes explanations of each slang term and the specific groups they typically reference. Team members could add newly discovered slang terms, which are then evaluated and annotated by the expert team. This dynamic maintenance ensures a consistent understanding of emerging language and slang among all team members.
Additionally, the document serves as a knowledge-sharing platform, providing the annotation team with consistent references and concrete examples for handling complex or ambiguous posts. This mechanism improves annotation efficiency and enhances the consistency of the annotations. The dynamically updated lexicon of annotated Chinese hateful slang become a valuable resource for research on Chinese hate speech detection.
3.4 Data Description
STATE ToxiCN dataset contains a total of 8,029 annotated posts, among which 4,942 posts include hateful content, accounting for 61.55%. A total of 9,533 quadruples are annotated, with 6,034 of them containing hateful information, making up 63.60%. We present the statistical details of STATE ToxiCN in Table 4. Gender, Region, and Race are the three most common group types in the dataset. Additionally, "multi-group" refers to Target-Argument pairs that involve hatred directed at multiple target groups, with a total of 854 such instances, accounting for 8.96%. In addition, the annotated lexicon includes 330 Chinese hateful slang terms collected from real online forums. More annotated examples can be found in Appendix A.
4 Experiment
4.1 Baselines
To evaluate the performance of LLMs with varying parameter sizes and capabilities in detecting span-level Chinese hate speech, we choose twelve well-known models across three categories:
Open-source Models: mT5-base Xue (2020), Mistral-7B Jiang et al. (2023), LLaMA3-8B AI@Meta (2024), Qwen2.5-7B Team (2024); Safety-domain Models: ShieldLM-14B-Qwen Zhang et al. (2024), ShieldGemma-9B Zeng et al. (2024), and Closed-source LLMs: LLaMA3-70B AI@Meta (2024), Qwen2.5-72B Team (2024), Gemini-1.5-Pro Team et al. (2023), Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), DeepSeek-v3 DeepSeek-AI et al. (2024).
4.2 Evaluation metrics.
We employ the Macro-F1 scores as the main evaluation metrics. The lack of word delimiters in Chinese introduces ambiguity in the extraction of targets and arguments. To obtain more accurate evaluation results, we therefore utilize both hard and soft matching metrics as our .
Hard-matching: A predicted quadruplet, particularly its target and argument components, is deemed correct only if it perfectly matches its corresponding gold label.
Soft-matching: We adopt the algorithm proposed by Han et al. (2023), where a prediction is considered correct if the Target and Argument scores achieve a threshold of 0.5.
Category #Posts Quad. Hateful Non-hate
Train 6424 7631 4842 2789
Test 1605 1902 1221 681
Total 8029 9533 6063 3470
Table 5: Statistics of train and test datasets, including #Posts, total quads (Quad.), hateful quads (Hateful), and non-hateful quads (Non-hate).
4.3 Experiment Settings
We conduct fine-tuning on open-source and safety-domain models, with training and testing set sizes detailed in Table 5. The fine-tuning process was performed using only a basic prompt, which included task definitions, output formats, and specific prediction requirements for all elements. We evaluate the performance of closed-source LLMs by calling their APIs. In addition to the aforementioned basic prompt, we also provided a hate speech example and a non-hate speech example. Further details regarding these requirements are provided in Appendix C.
5 Results and Analysis
Model Target Argument T-A Pair T-A-H Tri. Quad.
Hard Soft Hard Soft Hard Soft Hard Soft Hard Soft
Finetuned Models (with Basic Prompt)
mT5-base 59.15 70.55 28.63 67.03 23.33 55.90 17.76 43.34 16.60 38.61
Mistral-7B 62.97 73.69 35.58 70.90 30.55 60.49 26.15 51.01 23.72 45.62
LLaMA3-8B 64.07 73.74 36.72 70.82 31.64 60.88 27.04 51.62 24.27 46.08
Qwen2.5-7B 63.96 74.64 35.42 70.36 30.63 60.52 26.51 52.86 23.70 47.03
[…]
ShieldGemma-9B 63.40 74.31 34.40 71.11 29.99 61.51 25.64 52.70 23.49 47.14
LLM APIs (with Basic Prompt and 2 Examples)
LLaMA3-70B 30.54 41.03 14.39 47.96 8.16 27.34 6.03 20.70 3.69 11.93
Qwen2.5-72B 40.94 50.44 21.10 56.36 15.66 39.49 12.48 30.92 8.74 20.29
Gemini-1.5-Pro 29.80 37.29 18.43 54.96 9.37 26.22 7.71 21.88 5.45 14.81
Claude-3.5-Sonnet 37.61 50.72 15.45 57.24 9.72 36.16 7.94 29.82 6.29 22.45
GPT-4o 46.85 58.19 22.64 62.41 17.21 46.41 13.21 35.68 9.00 23.34
DeepSeek-v3 48.16 59.25 22.79 59.38 18.68 46.40 14.95 37.19 11.48 27.38
Table 6: Performance comparison of various models across different levels of annotated tasks, including Target, Argument, Target-Argument Pair (T-A Pair), Target-Argument-Hateful Triple (T-A-H Tri.), and Target-Argument-Hateful-Group Quadruple (Quad.) under both Hard and Soft evaluation metrics.
5.1 RQ1: Can LLMs identify the spans of targets and arguments in Chinese text?
Finetuned Models
In the tasks of identifying targets, arguments, and target-argument pairs (T-A Pair), fine-tuned models significantly outperform direct usage of LLM APIs. LLaMA3-8B, Qwen2.5-7B, and the Shield series models all achieve hard match scores of over 63%, strongly demonstrating the superior ability of fine-tuned models in identifying text span boundaries. Soft match metrics further confirm the advantage of fine-tuned models in target and argument identification, with F1 score approaching 70% or higher. These results indicate that fine-tuned models on task-specific data show a substantial advantage in span identification. LLaMA3-8B and Qwen2.5-7B perform the best. ShieldGemma-9B achieves the highest scores in soft-match metrics for two tasks.
LLM APIs
Compared to fine-tuned models, LLMs access through APIs performed significantly worse in all tasks, regardless of whether hard or soft match metrics were used. Even with few-shot prompting strategies, their performance still lag far behind that of fine-tuned models. This suggests that LLMs, without being fine-tuned on Chinese hate speech data, are not adept at identifying text span boundaries, which is consistent with the fact that their training objectives are not directly related to this task. Specifically, the hard match scores for LLMs on "Target" and "Argument" were only between 40-50%, while soft match scores were around 60%. GPT-4o and DeepSeek-v3 exhibit the best performance.
Comparison and Summary
The fine-tuning technology can improve models’ abilities in identifying text span boundaries, demonstrating a clear advantage in target and argument extraction tasks. Joint extraction tasks are more challenging than single element extraction tasks, and all models perform worse in argument extraction than in target extraction. We believe this is primarily because in Chinese text, argument spans tend to be longer and structurally more complex, leading to poor hard match performance. However, the soft match scores for argument extraction were similar to those of target extraction, indicating that the models’ semantic understanding capabilities are comparable in both tasks. This suggests that despite difficulties in identifying precise boundaries, the models remain effective in semantic understanding.
Model Target Argument T-A Pair T-A-H Tri. Quad.
Hard Soft Hard Soft Hard Soft Hard Soft Hard Soft
Finetuned Models (with Basic Prompt)
mT5-base 56.832.32 68.332.22 27.171.46 64.172.86 21.332.00 51.174.73 18.171.46 44.671.33 16.330.27 36.172.44
Mistral-7B 61.031.94 72.621.07 36.711.13 70.050.85 30.270.28 58.621.87 26.250.10 51.050.04 22.221.50 43.002.62
LLaMA3-8B 61.262.81 72.121.62 35.171.55 70.830.01 28.533.11 59.001.88 24.472.57 51.380.24 19.774.50 42.463.62
Qwen2.5-7B 61.792.17 73.331.31 36.260.82 69.590.77 29.920.71 57.722.80 27.641.13 53.660.80 22.760.94 45.041.99
[…]
ShieldGemma-9B 62.500.90 74.350.04 35.711.31 71.100.01 29.870.12 60.710.80 26.951.31 55.032.33 23.210.28 46.101.04
LLM APIs (with Basic Prompt and 2 Examples)
LLaMA3-70B 30.870.33 41.450.42 14.800.41 46.681.28 8.290.13 25.381.96 7.401.37 22.581.88 4.721.03 13.141.21
Qwen2.5-72B 45.584.64 54.674.23 22.151.05 57.361.00 16.771.11 41.612.12 12.420.06 30.350.57 8.830.09 19.590.70
Gemini-1.5-Pro 31.521.72 39.071.78 18.940.51 54.570.39 9.010.36 27.951.73 8.610.90 25.833.95 6.230.78 17.352.54
Claude-3.5-Sonnet 41.453.84 54.063.34 15.800.35 55.801.44 10.430.71 36.960.80 9.281.34 33.043.22 7.100.81 25.222.77
GPT-4o 49.282.43 61.303.11 21.710.93 59.662.75 16.520.69 46.550.14 12.011.20 33.721.96 8.460.54 22.800.54
DeepSeek-v3 52.694.53 64.254.00 23.791.00 62.102.72 19.220.54 50.404.00 16.131.18 42.074.88 11.690.21 30.112.73
Table 7: Performance comparison of various models on a hateful-slang-containing subset of the test set. Green indicates a decrease in F1 score on the subset compared to the full dataset, while red indicates an increase.
5.2 RQ2: Can LLMs classify Chinese target-argument pairs as hateful and identify the groups?
Finetuned Models
Fine-tuned models demonstrate a clear advantage in both the triplet and quadruplet tasks. In the Target-Argument-Hateful triplet task (T-A-H Tri.), the fine-tuned models achieve hard-match metrics of around 26% and soft-match metrics of around 52%. In the more complex quadruplet task, the hard-match metrics of the fine-tuned models are mostly between 25% and 27% when simultaneously identifying hateful content and target groups, while the soft-match metrics are approximately 45% - 47%. This indicates that the fine-tuned models’ performance is mainly limited by the target-argument pair identification. Overall, although the fine-tuned models perform poorly in hard-match metrics, the results in soft-match metrics reach approximately 50%. Among these, LLaMA-3-8B performs best in hard-match, while Qwen2.5-7B and ShieldGemma-9B perform better in soft-match metrics.
LLM APIs
LLM APIs perform poorly on both the triplet and quadruplet tasks. In the T-A-H triplet task, LLMs achieve hard-match metrics of only 6% - 15% and soft-match metrics of 20% - 38%. In the quadruplet task, the hard-match metrics of LLMs further drop to 3% - 12%, and the soft-match metrics, though slightly better, are still only 11% - 27%. We believe that the performance limitations also primarily stem from the difficulty in identifying accurate span boundaries. These data clearly show that LLMs, without being fine-tuned on Chinese hate speech data, are significantly inadequate at performing Chinese hate speech triplet and quadruplet predictions, even with few-shot prompting failing to bridge the performance gap. DeepSeek-v3 performs significantly better than other APIs, particularly in the quadruplet prediction task.
Comparison and Summary
The fine-tuned model significantly outperforms the LLM APIs, with the performance gap primarily stemming from T-A pair extraction. The fine-tuned model outperforms the LLM APIs in both hate speech the triplet and quadruplet prediction. Specifically, the fine-tuned models show better performance in tasks requiring the simultaneous identification of hate speech and its target group. Although there is still room for improvement in terms of precise matching, the application of fine-tuning technology undoubtedly provides a feasible path for handling such complex tasks. At the same time, the performance of the LLM APIs in these tasks is disappointing, indicating that they need targeted fine-tuning to be applied to these types of scenarios.
Figure 1: Examples of Chinese hateful slang understanding analysis with LLM. The hateful slang terms and texts appear in black, ShieldGemma-v3 explanations are in green, DeepSeek-v3 explanations are in red, and human annotator explanations are in blue.
5.3 RQ3: Can LLMs understand Chinese hateful slang?
5.3.1 Impact of Hateful Slang on Chinese Hate Speech Detection
To evaluate the impact of hate slang on span-level Chinese hate speech detection models, we select 502 posts containing hate slurs from the test set for focused testing. By comparing the model performance on this subset of data with its performance on the full test set, we can gain a clearer understanding of how hate slurs affect model performance. The experimental results are presented in Table 7. The superscript numbers in the table indicate the difference in F1 score between the model’s performance on the selected subset and its performance on the full test set. Green indicates a decrease in F1 score on the subset compared to the full dataset, while red indicates an increase.
Finetuned Models
Across most tasks, the hard and soft matching metrics for the models show a decreasing trend. However, the metric for the hate speech triplet prediction task (T-A-H Tri.) is an exception, exhibiting an increase. We attribute this primarily to two factors: Firstly, posts containing hateful slang generally demonstrate a more explicit hateful intent, while the posts we remove include many types that are difficult to identify, such as implicit hate expressions. This makes it easier for the T-A-H Tri. task to capture the relationships within hate speech triplets, thereby leading to the observed increase in the metric. Secondly, the experimental results also confirm that the main challenges posed by hateful slang lie in extracting the target and argument, as well as accurately classifying the targeted group. When the training data is not sufficiently comprehensive, fine-tuned models struggle to effectively handle the complexities introduced by hateful slang.
LLM APIs
Surprisingly, LLM APIs demonstrate superior performance on the subset of the test set containing hateful slang. With the exception of a slight decrease in the soft matching metric for argument identification, nearly all other metrics show an improvement. We believe that this is primarily due to two factors: First, hateful slang often appears as the target, and the explicit target terms effectively compensate for the limitations of LLM APIs in span boundary recognition. Second, the identification of hateful slang often requires rich background knowledge; for example, Example 2 in the figure requires knowledge of Chinese history and related common usages, which is precisely where LLM APIs excel. Therefore, hateful slang, which is difficult for fine-tuned models to leverage, becomes an effective source of information for LLM APIs.
Comparison and Summary
Fine-tuned models and LLM APIs exhibit distinct characteristics when processing hateful slang. Fine-tuned models, when lacking sufficient training and understanding of specific domain knowledge, struggle to effectively address the challenges posed by hateful slang, particularly in areas such as fine-grained entity recognition and complex language pattern understanding. In contrast, LLM APIs leverage their robust background knowledge and contextual understanding to better comprehend and handle these complex situations, thus achieving superior performance. However, the performance gap between LLM APIs and fine-tuned models on this task remains significant. Infusing the background knowledge of large models into fine-tuned models may offer a promising avenue for improvement.
5.3.2 Understanding Capabilities of LLMs on Chinese Hateful Slang
To investigate the understanding capabilities of large language models regarding Chinese hateful slang, we conducted experiments using two relatively high-performing models, ShieldGemma-9B and DeepSeek-v3. In these experiments, we provided the models with hateful terms and their surrounding sentence context, asking them to generate explanations of the hateful slang and identify the potentially affected groups. The experimental results are detailed in Figure 1, with specific experimental details provided in Appendix B.
For a more nuanced analysis of the models’ understanding abilities, we selected two terms with distinct Chinese characteristics as case studies. In Case 1, "默" (mò) is a typical "merging word," whose literal meaning is ’silence,’ but it is composed of the characters "黑犬" (hēiquǎn), meaning "black dog." This type of character combination is a unique linguistic phenomenon in Chinese. The experimental results showed that neither model could accurately understand the hateful information contained within it. In Case 2, "冉闵" (Rǎn Mǐn) is a hateful slang term rooted in Chinese history and culture, referring to an ancient emperor who massacred ethnic minorities and is often used for racial discrimination. In comparison, while ShieldGemma-9B could understand the background information, it failed to generate an accurate explanation of the hateful information and clearly identify the potentially affected groups. DeepSeek-v3’s generated information, however, closely matched the results given by human annotators. This demonstrates that even relatively high-performing models face significant challenges in understanding Chinese hateful slang that possesses cultural specificity and subtle connotations.
6 Conclusions
Different from previous work focusing on template design, we address implicit sentiment detection from a human cognitive perspective. In this paper, we propose a straightforward and effective strategy-level approach inspired by dual-process theory, STATE ToxiCN, which detects implicit sentiment by a dual-system reasoning framework in the ASQP task. Firstly, we employ commonsense reasoning tools to estimate human reactions and form intuitive reaction augmentation samples. Secondly, we integrate the augmented text with analytic system-based instruction and use this information to enable the model to further integrate contextual understanding and intuitive reaction augmentation samples, resulting in a comprehensive construction of the sentiment quads. The experimental results show that our STATE ToxiCN can achieve state-of-the-art performance compared to existing baseline methods. We also conduct experiments in various scenarios to thoroughly demonstrate the effectiveness, universality, and robustness of STATE ToxiCN.
Limitation
Despite implementing rigorous quality control measures in constructing the span-level Chinese hate speech detection dataset and the annotated lexicon, we acknowledge several limitations. First, although we have taken steps to minimize labeling bias, subjective differences in annotators’ understanding of toxic language may still result in mislabeled data. Additionally, while we strived to ensure accuracy and consistency in annotating target-argument pairs, the inherent characteristics of the Chinese language, such as flexible grammar and ambiguous boundaries, pose challenges for completely precise annotations.
Our Annotated Lexicon covers a wide range of commonly used Chinese hateful slang; however, due to the dynamic nature and rapid evolution of the Chinese internet language, certain emerging or less widely recognized cloaking expressions (e.g., more complex word transformations or domain-specific slang) might not be fully captured. Finally, this study focuses primarily on textual features and does not consider non-textual elements, such as images, videos, or metadata about the authors, which may limit the model’s ability to comprehensively detect multimodal hate speech.
In the future, we aim to expand the scope of our dataset to include more diverse contexts and cloaking language, while exploring multimodal and cross-domain hate speech detection methods to improve overall performance and applicability.
Ethics Statement
We adhere strictly to the data usage agreements of all public online social platforms and conduct thorough reviews to ensure that no user privacy information is included in our dataset. The opinions and findings reflected in the samples of our dataset do not represent the views of the authors, either explicitly or implicitly. We aim to ensure that the benefits of our proposed resources outweigh any potential risks. All resources are provided exclusively for scientific research purposes.
Appendix A Details of the Annotated Lexicon
We use JSON files to store the annotated lexicon, which is a mainstream coding specification to facilitate machine-readability. The data structure is defined as Sample = (Term, [Groups], Definition), where Groups indicates the groups that are commonly targeted by the hateful slang. If a term does not have a specific targeted group, then Groups is set to "others". Here we provide four samples in Figure 2.
Figure 2: Examples of annotated Chinese hateful slang, including Term, Groups and Definition.
Appendix B Experimental Details of Hateful Slang Understanding Generation
We prompt the non-fine-tuned ShieldGemma-9B and DeepSeek-v3 to generate hateful slang understanding. The prompt used is as follows.
Prompt: [Text]中的[Term]是否包含特定群体的仇恨信息,如果是,请说明对哪些群体构成仇恨并说明原因。
Prompt in English: In [Text], does the [Term] constitute hate speech directed at specific groups? If yes, please identify the targeted groups and the reasons behind the hatred.
Appendix C Details of the Prompt
The basic prompt and the version with 2 examples are shown in Figure 3.
Figure 3: Illustration of the basic prompt and the version with 2 examples.